### v3

@@ Abstract
Image-based virtual try-on aims to synthesize a naturally dressed person image with a clothing image, which revolutionizes online shopping and inspires related topics within image generation, showing both research significance and commercial potential. However, there is a gap between current research progress and commercial applications and an absence of comprehensive overview of this field to accelerate the development. In this survey, we provide a comprehensive analysis of the state-of-the-art techniques and methodologies in aspects of pipeline architecture, person representation and key modules such as try-on indication, clothing warping and try-on stage. We propose a new semantic criteria with CLIP, and evaluate representative methods with uniformly implemented evaluation metrics on the same dataset. In addition to quantitative and qualitative evaluation of current open-source methods, unresolved issues are highlighted and future research directions are prospected to identify key trends and inspire further exploration. The uniformly implemented evaluation metrics, dataset and collected methods will be made public available at https://github.com/little-misfit/Survey-Of-Virtual-Try-On.
@@ 1 Introduction
To overcome the above difficulties, tremendous efforts have been made and Fig. 1 show some representative methods on a timeline. In 2017, CAGAN [50] gave the first try by employing CycleGAN [141] to overcome the lack of training triplet data, i.e., (original person image, in-shop clothing image, try-on image), but the generation quality is far from satisfactory. Subsequently, VITON [37] creatively proposed clothing-agnostic person representation by human parsing to make up the lack of supervised training data. They constructed the basic network framework of “Try-On Indication + Cloth Warping + Try-on”, laying the foundation for further improvement on generation quality in subsequent works [82, 119, 75, 111, 130, 3, 4, 60, 135, 61]. Apart from continuous quality improvements, some new goals are desired. Trying on multiple garments attracted attention in 2019 [58], but such works are scarce due to the complex interaction between multiple garments and the lack of specialized datasets [85, 65, 22]. In order to remove the heavy reliance on human parsing, several teacher-student networks were designed [47, 31, 40] to achieve parser-free try-on at inference time. Recently, with the advancement of computational capabilities, high-resolution virtual try-on tasks has become possible, e.g., VITON-HD [18] and HR-VTON [61]. Inspired by StyleGAN [54, 55, 73] and Diffusion model [43, 87, 112, 70] in the field of image generation, single-stage networks [62, 27] as well as Diffusion architecture [142, 56, 136] emerged for this task.
@@
  • Evaluation: unified evaluation, new criterion and user study. We evaluate open-source works with the same dataset, and also perform user study on visual results with 333 volunteers. Particularly, we compute Semantic Score with CLIP [94] as a new criterion which can respectively evaluate the semantic similarity of try-on and non-try-on parts. Relevant data and codes will be publicly available at https://github.com/little-misfit/Survey-Of-Virtual-Try-On.
@@
Table 1: Key characteristics of representative image-based virtual try-On methods. Please refer the number of pipeline to Fig. 2 and the number of person representation is shown in Fig. 4. The three loss items are the respective losses for modules of Try-On Indication, Cloth Warping and Try-on. “-” means this item is not available or not indicated in the published work.
@@ 2.1 Overview
It is worth noting that the above three steps are not necessarily present at the same time, and there is no strict order. Table 1 summarizes representative methods and we will discuss the key designs in the following subsections. The statistics of image-based virtual try-on methods are summarized from the literature listed in Table 1 and the following methods [3, 4, 2, 45, 64, 39, 103, 98, 28, 80, 93, 104, 49, 95, 44, 113, 24, 19, 88, 14, 127, 76, 140, 53, 59, 91, 125, 123, 63, 72, 121, 124, 29, 6, 92, 134, 139, 71, 86, 90, 106, 1].
@@ 2.5.3 Flow Estimation
Dior [22] employs a single-layer flownet to deform the clothing features, where GFFE (Global Flow Field Estimator) proposed by [99] is used to predict flow map and soft mask. To capture more 3D changes of human posture deformation, 3D-GCL [46] uses flownet [68] to create SMPL [78] flows as ground truth to constrain the flow map, which results in better visual outcomes on the Deep Fashion dataset.
@@ 2.5.4 Implicit Transformation
As shown in Fig. 5(d), there has been a diversification in clothing warping methods in recent years. Single-STN methods are unable to handle natural deformation. Apart from a few works employing Multi-STN, most methods use a combination of STN and TPS, and they are classified under the TPS category. Influenced by the GMM framework, TPS transformation is the most popular method. TPS transformations have limitations in terms of degrees of freedom for deformation. In order to achieve more diverse clothing warping results, flow estimation has become a breakthrough in deformation performance. At the cost of greater computational overhead, flownet has achieved stronger deformation capabilities, and its usage is on the rise. Implicit transformation methods deform garments in a generative manner, further enhancing the diversity of garment deformation but also posing challenges in controlling clothing content. With the development of image generation technology under the diffusion framework, implicit transformation methods become promising, whereas diffusion based methods usually directly synthesize the final try-on image without explicitly warping the clothes in advance.
@@ 2.6 Try-On
Table 2: Key properties of datasets. Popularity refers to the number of usage that is summarized from existing methods.
@@
DeepFashion 2016
$1101\times 750$
52,712/-
11
VITON-HD 2021
$1024\times 768$
11,647/2,032
15
ESF 2022
@@
Dress Code 2022
$1024\times 768$
48,392/5,400
11
Previously, only using the Try-On module (without explicit try-on indication and cloth warping) could only generate rough try-on images. However, this issue has been alleviated in the era of large-scale models. TryOnDiffusion [142] has designed a substantial Parallel-UNet, which, after being trained with massive amounts of data, can directly use the original image and pose features to accomplish high-quality try-on. More recent diffusion based methods put additional emphasis on attention mechanism [56] and acceleration strategy [136].
@@
3 Datasets and Evaluation Criteria
@@ 3.1 Datasets
In this section, we will introduce the currently public virtual try-on datasets. Before showing the details of each dataset, Table 2 compares them in terms of several key characteristics and counts the usage quantities of various datasets. VITON [37] is the first dataset for image-based virtual try-on and also the most popular one. Fig. 6 visually shows the samples for each dataset.
VITON dataset was released by [37], which collected a total of 16,253 pairs of top clothing images and frontal-view woman images wearing corresponding clothes. The clothing images were taken from the front with the clothing in a flat state and the background was pure white. The person body images were taken from the front upper body of the model, and the background was gray or light gray. 14,221 pairs of images are used as the train set, and 2,032 pairs of images are used as the test set.
Figure 6: Examples in the dataset. VITON contains pairs of female photo and corresponding in-shop clothing image. MPV cosists of triplets of female photos in different poses with the same clothing and the corresponding clothing image. It also parses pose keypoints. Deep Fashion (In-Shop Clothes Retrieval Benchmark) contains images of the same model in different poses. VITON-HD is a high-resolution dataset accompanied with parsed segmentation, dense pose and pose keypoints. ESF consists of clothed model photos for both male and female. Dress Code involves more clothing types and provides dressed model photo, clothing image, pose keypoints, segmentation and dense pose.
MPV dataset was released by [23], which contains 35,687 model images and 13,524 clothing images. The clothing images are the front images of the clothing in the flat state, and the model images are light gray background images. Each model in MPV corresponds to multiple images of different poses, and the image resolution is $256\times 192$. Dong et al. extracted 62,780 triplets of the same person wearing the same clothes but with different poses, further dividing 52,236 triplets as the training set, and shuffling 10,544 triplets as the test set.
Deep Fashion [77] is a large-scale dataset released by the Chinese University of Hong Kong. It contains 800,000 images, including different angles, scenes, buyer shows, seller shows, etc. There are four subsets in total, namely: Category and Attribute Prediction Benchmark, In-Shop Clothes Retrieval Benchmark, Consumer-to-shop Clothes Retrieval Benchmark, and Fashion Landmark Detection Benchmark. The In-Shop Clothes Retrieval Benchmark is often used for virtual try-on tasks, which contains product IDs, images of the same model in different poses, with a resolution of $1101\times 750$. There are a total of 7,982 products corresponding to 52,712 images. Each image is annotated with a bounding box, clothing type, and pose type. Most of the clothes in this dataset are put on bodies with rare separate clothing images.
VITON-HD dataset was released by [18], with a resolution of $1024\times 768$. As shown in Fig. 6, it additionally contains parsing information such as segmentation, dense pose and pose keypoints. VITON-HD is divided into 11,647 image pairs for training and 2,032 image pairs for testing.
E-Shop Fashion(ESF) dataset was proposed by [27], which contains 180,000 clothed model photos with adjusted resolution of $512\times 512$. Each image contains a front-standing pose model, and the model image is centrally aligned and filled to the region between the chin and thighs. Furthermore, the dataset is split into a training set containing 170,000 samples and a test set containing 10,000 samples.
Dress Code dataset was proposed by [83], which is more than three times larger than the publicly available virtual try-on datasets, with high-resolution paired images ($1024\times 768$). Different from VITON, which only contains upper-body female image pairs, Dress Code contains upper-body, lower-body, and full-body clothing for both female and male, as well as corresponding model try-on images. The images are divided into three categories: upper-body clothing (consisting of tops, T-shirts, shirts, sweatshirts and sweaters), lower-body clothing (consisting of skirts, pants, shorts, and leggings), and dresses. In total, the dataset consists of 53,795 image pairs: 15,366 pairs for upper-body clothing, 8,951 pairs for lower-body clothing, and 29,478 pairs for dresses. It is divided into a training set containing 48,392 pairs and a test set containing 5,400 pairs (1,800 pairs for upper body, lower body, and full body, respectively). The images are pre-processed and cut at the nose to ensure the privacy of the model. Dress Code also provides pose keypoints, dense pose, and segmentation mask for clothed model images.
In summary, although VITON is the most popular dataset, its limitations include a single type of clothing and a small amount of data. As a ground breaking work for image generation based virtual try-on, it could be used as the most basic benchmark with most works for comparison. However, it restricts the generalization performance. The most recently published Dress Code is a comprehensive dataset with the most diverse types of in-shop clothing images and the widest variety of models for trying on clothes, which could be a good choice for training virtual try-on models at present.
@@ 3.2 Evaluation Criteria
SSIM: Structural Similarity Index Metric [122] is a traditional image similarity evaluation standard that is originally used for evaluating the performance of image compression algorithms. SSIM index is influenced by the theory of human visual system (HSV) and is calculated from the three aspects of image brightness, contrast and structure. The inputs to compute SSIM are two individual images, i.e., the ground truth image and the generated image. SSIM value ranges within [0,1], where the larger value indicates the smaller difference between the compared two images and the better realism of the generated image. This criterion is sensitive to pixel shift, so high-level semantic feature is also considered for the evaluation.
@@
IS: Inception Score [105] is a metric based on Inception network, which is originally used to evaluate generation models in terms of clarity and diversity. The calculation of IS only involves generated images, and the prediction probability of generated images is used to describe the performance. Clarity refers to more certain probability of belonging to one class, e.g., the prediction entropy is smaller. Diversity involves a set of generated images, where the predictions are evenly distributed across all classes indicate better diversity. The higher value of IS indicate the better generation performance. However, different from the original generation tasks, the generation target for virtual try-on is a fixed class such as human photos. It is not fair enough for this task (therefore not adopted in our unified evaluation), and FID becomes an alternative way which compares the distributions of generated images and ground truth.
@@
FID: Frechet Inception Distance [42] measures the statistical similarity between two sets of images by calculating the Frechet distance between the feature vectors of the real images and the generated images. Lower value of FID indicate smaller difference between two sets in feature space and better performance.
@@
LPIPS: Like IS and FID, Learned Perceptual Image Patch Similarity [137] also utilizes a pretrained network (e.g., VGG [108] and AlexNet [57]) to capture high-level features. It calculates Euclidean or cosine distance between features output via multiple layers. Lower value of LPIPS indicate closer similarity of two images.
Semantic Score: We propose a new semantic score with the help of recent advanced model CLIP [94] which connects text and image. The Euclidean distance between CLIP features of generated image and ground truth is calculated, so lower value indicate closer similarity in semantic information.
@@
It is worth noting that we for the first time separately evaluate the clothing warping performance and generation ability for non-try-on area with a human parser [67]. It further shows which module contributes more to existing methods.
Figure 7: Quantitative results on VITON. The vertical axis indicates the testing condition is paired or unpaired and the evaluation target is clothing, non-try-on area (denoted as “person”) or all together. For method symbol, the cross symbol indicates the clothing warping strategy is flow while the triangle symbol represents TPS-based method. *Both HR-VTON and VITON-HD are tested using high-resolution ($1024\times 768$) datasets. To compare with other low-resolution models, we down-sample the high-resolution results to low resolution ($256\times 192$). To illustrate the impact of image resolution on evaluation metrics, we denote the scores of high-resolution results with gray points.

### v4

@@ Abstract
Image-based virtual try-on aims to synthesize a naturally dressed person image with a clothing image, which revolutionizes online shopping and inspires related topics within image generation, showing both research significance and commercial potential. However, there is a gap between current research progress and commercial applications and an absence of comprehensive overview of this field to accelerate the development. In this survey, we provide a comprehensive analysis of the state-of-the-art techniques and methodologies in aspects of pipeline architecture, person representation and key modules such as try-on indication, clothing warping and try-on stage. We additionally apply CLIP to assess the semantic alignment of try-on results, and evaluate representative methods with uniformly implemented evaluation metrics on the same dataset. In addition to quantitative and qualitative evaluation of current open-source methods, unresolved issues are highlighted and future research directions are prospected to identify key trends and inspire further exploration. The uniformly implemented evaluation metrics, dataset and collected methods will be made public available at https://github.com/little-misfit/Survey-Of-Virtual-Try-On.
@@ 1 Introduction
To overcome the above difficulties, tremendous efforts have been made and Fig. 1 show some representative methods on a timeline. In 2017, CAGAN [51] gave the first try by employing CycleGAN [144] to overcome the lack of training triplet data, i.e., (original person image, in-shop clothing image, try-on image), but the generation quality is far from satisfactory. Subsequently, VITON [38] creatively proposed clothing-agnostic person representation by human parsing to make up the lack of supervised training data. They constructed the basic network framework of “Try-On Indication + Cloth Warping + Try-on”, laying the foundation for further improvement on generation quality in subsequent works [85, 123, 78, 115, 133, 3, 4, 61, 138, 62]. Recently, with the advancement of computational capabilities, high-resolution virtual try-on tasks has become possible, e.g., VITON-HD [17] and HR-VTON [62]. Inspired by StyleGAN [55, 56, 76] and Diffusion model [44, 90, 116, 73] in the field of image generation, single-stage networks [63, 27] as well as Diffusion architecture [145, 57, 139] emerged for this task.
Apart from continuous quality improvements, some new goals are desired. In order to remove the heavy reliance on human parsing, several teacher-student networks were designed [48, 32, 41] to achieve parser-free try-on at inference time. Paired training requires tremendous paired data and does not fit well in the unpaired try-on scenario. Previous efforts such as interpolation in latent space [63], uv mapping via dense pose [22] and disentanglement of garment style and spatial features [127, 129] have been made, to facilitate unpaired image-based virtual try-on. Trying on multiple garments, that is, mix-match, has practical values but is challenging in dealing with the interaction between garments (e.g., tuck or untuck). Generating semantic layout [88, 66] or giving the dressing order [21, 129] for different garments are two dominant solutions. Controllable try-on in regards of clothing sizes [15] and types [132, 68] are achieved via associated skeleton or landmarks. With the development of mask inpainting in the image generation area, recently in-the-wild try-on becomes popular [69, 22, 91, 57], where the background is more natural and complex.
@@
  • Evaluation: unified evaluation, new criterion, and user study. We evaluate open-source works with the same dataset, and also perform a user study on visual results with 139 volunteers. Particularly, we separately evaluate try-on and non-try-on parts and additionally apply CLIP [98] to measure the semantic similarity. Relevant data and codes will be publicly available at https://github.com/little-misfit/Survey-Of-Virtual-Try-On.
@@
Table 1: Key characteristics of representative image-based virtual try-On methods. Please refer the number of pipeline to Fig. 2 and the number of person representation is shown in Fig. 4. The three loss items are the respective losses for modules of Try-On Indication, Cloth Warping and Try-on.
@@
UniHuman
[69]
CVPR \@slowromancapi@
1,5,9,13
- - - - Diffusion
LH-400K
WPose
PICTURE
[91]
CVPR \@slowromancapiii@
4,8,12
Diffusion - - - Diffusion
WUTON
DeepFashionM
SHHQ
VITON-HD
@@ 2.1 Overview
It is worth noting that the above three steps are not necessarily present at the same time, and there is no strict order. Table 1 summarizes representative methods and we will discuss the key designs in the following subsections. The statistics of image-based virtual try-on methods are summarized from the literature listed in Table 1 and the following methods [3, 4, 2, 46, 65, 40, 107, 102, 28, 83, 97, 108, 50, 99, 45, 117, 24, 18, 92, 14, 131, 79, 143, 54, 60, 95, 129, 64, 75, 125, 128, 30, 6, 96, 137, 142, 74, 89, 94, 110, 1].
@@ 2.5.3 Flow Estimation
Dior [21] employs a single-layer flownet to deform the clothing features, where GFFE (Global Flow Field Estimator) proposed by [103] is used to predict flow map and soft mask. To capture more 3D changes of human posture deformation, 3D-GCL [47] uses flownet [71] to create SMPL [81] flows as ground truth to constrain the flow map, which results in better visual outcomes on the Deep Fashion dataset. RP-VTON [68] utilizes landmarks to control cloth warping, which is novel in achieving controllable try-on.
@@ 2.5.4 Implicit Transformation
As shown in Fig. 5(d), there has been a diversification in clothing warping methods in recent years. Single-STN methods are unable to handle natural deformation. Apart from a few works employing Multi-STN, most methods use a combination of STN and TPS, and they are classified under the TPS category. Influenced by the GMM framework, TPS transformation is the most popular method. TPS transformations have limitations in terms of degrees of freedom for deformation. In order to achieve more diverse clothing warping results, flow estimation has become a breakthrough in deformation performance. At the cost of greater computational overhead, flownet has achieved stronger deformation capabilities, and its usage is on the rise. Implicit transformation methods deform garments in a generative manner, further enhancing the diversity of garment deformation but also posing challenges in controlling clothing content. With the development of image generation technology under the diffusion framework, implicit transformation methods become promising.
@@ 2.6 Try-On
Table 2: Key properties of datasets. Popularity refers to the number of usage from existing methods.
@@
SHHQ 2022 S
$1024\times 512$
231,176/-
2
UPT 2021 S
$512\times 320$
27,139/6,115
2
ESF 2022 S
@@
StreetTryOn 2024 W
$512\times 320$
12,364/2,089
1
LH-400K 2024 W
$512\times 512$
409,270/-
1
WPose 2024 W
-/2,304
1
WVTON 2024 W
-/440
1
Previously, only using the Try-On module (without explicit try-on indication and cloth warping) could only generate rough try-on images. However, this issue has been alleviated in the era of diffusion models. TryOnDiffusion [145] has designed a substantial Parallel-UNet, which, after being trained with massive amounts of data, can directly use the original image and pose features to accomplish high-quality try-on. More recent diffusion-based methods put additional emphasis on attention mechanism [57], acceleration strategy [139], and image editing techniques [91, 69].
@@
Figure 6: Examples in existing datasets.
3 Datasets and Evaluation Details
@@ 3.1 Datasets
The key properties of existing datasets for image-based virtual try-on are illustrated in Table 2, and Fig. 6 show some examples within these datasets. The datasets containing clothing images are commonly used for trying on in-shop clothes with paired annotation while those without clothing images usually demonstrate the performance of unpaired try-on. The property of multi-pose in Table 2 does not refer to the pose types within the whole dataset but indicates whether one individual sample owns multiple poses.
VITON [38] was the first collection for the task of trying on in-shop clothes and opens up a new research area, but currently are not publicly available due to ethical concerns. VITON-HD [17] is an alternative dataset, with a resolution of $1024\times 768$. It additionally contains parsing information such as segmentation, dense pose and pose keypoints. Deep Fashion [80] is a large-scale dataset including four subsets namely: Category and Attribute Prediction Benchmark, In-Shop Clothes Retrieval Benchmark, Consumer-to-shop Clothes Retrieval Benchmark, and Fashion Landmark Detection Benchmark. The In-Shop Clothes Retrieval Benchmark is often used for virtual try-on tasks, which contains product IDs and images of the same model in different poses. Each model in MPV [23] corresponds to multiple images of different poses. Dress Code [86] contains upper-body, lower-body, and full-body clothing for both female and male, as well as corresponding model try-on images which are cut at the nose to ensure the privacy.
SHHQ [29] aims at full-body human generation and has features of diverse attributes such as human poses, and characters’ garment texture and type. UPT [127] contains person images in half-body and full-body frontal views, wearing a variety of garments. The clothed model images in E-Shop Fashion (ESF) dataset [27] are centrally aligned and filled to the region between the chin and thighs.
Recently, in-the-wild try-on attracts interests and the datasets where person images have diverse backgrounds are collected. StreetTryOn [22] is a new benchmark for evaluating virtual try-on in natural environments with a wide range of clothing categories and human poses. LH-400K, WPose and WVTON are proposed by [69] with a variety of backrgounds, ages and body sizes, respectively serving as training and testing datasets. Specifically, WPose focuses on diverse poses and WVTON concentrates on garments.
@@ 3.2 Evaluation Criteria
SSIM: Structural Similarity Index Metric [126] is a traditional image similarity evaluation standard that is originally used for evaluating the performance of image compression algorithms. The inputs to compute SSIM are two individual images, i.e., the ground truth image and the generated image. SSIM value ranges within [0,1], where the larger value indicates the smaller difference between the compared two images and the better realism of the generated image. This criterion is sensitive to pixel shift, so high-level semantic feature is also considered for the evaluation.
@@
IS: Inception Score [109] is a metric based on Inception network, which is originally used to evaluate generation models in terms of clarity and diversity. The calculation of IS only involves generated images, and the prediction probability of generated images is used to describe the performance. The higher value of IS indicates the better generation performance. However, different from the original generation tasks, the generation target for virtual try-on is a fixed class such as human photos. It is not fair enough for this task (therefore not adopted in our unified evaluation), and FID becomes an alternative way which compares the distributions of generated images and ground truth.
@@
FID: Frechet Inception Distance [43] measures the statistical similarity between two sets of images by calculating the Frechet distance between the feature vectors of the real images and the generated images. Lower value of FID indicates smaller difference between two sets in feature space and better performance.
@@
LPIPS: Like IS and FID, Learned Perceptual Image Patch Similarity [140] also utilizes a pretrained network (e.g., VGG [112] and AlexNet [58]) to capture high-level features. It calculates Euclidean or cosine distance between features output via multiple layers. Lower value of LPIPS indicates closer similarity of two images.
Semantic Score: We propose a new semantic score with the help of recent advanced model CLIP [98] for evaluating image-based virtual try-on. The Euclidean distance between CLIP features of generated image and ground truth is calculated, so lower value indicate closer similarity in semantic information. Compared with previous pretrained network such as VGG and AlexNet used in IS/FID/LPIPS, CLIP is suited for this task due to its training on a diverse range of images and text, enabling it to capture a broad spectrum of visual and semantic features. Unlike traditional metrics like SSIM that primarily focus on pixel-level accuracy, CLIP assesses the semantic coherence between the generated and ground truth images, which is critical in evaluating how realistically the garments are portrayed in virtual try-on. Fig. 7 also shows that the semantic score has advantages in evaluating the semantic difference for virtual try-on. The scores above and below the subfigure are computed by comparing the generated result with ground truth (i.e., the left-most subfigure). Although both LPIPS and CLIP indicate that the right result is better, CLIP exhibit a large gap between the two results.
@@
It is worth noting that we for the first time separately evaluate the clothing warping performance and generation ability for non-try-on area with a human parser [70]. The generation quality of try-on area and non-try-on area reflects the fidelity of clothing warping and the preservation of human characteristics, which could provide existing methods valuable information for further improvement.
Figure 7: Comparison between semantic score and LPIPS. The scores for different generated results indicate that the proposed semantic score (denoted as CLIP in the figure) is more sensitive to semantic difference.
3.3 Implementation Details
Due to the ethical concern raised by VITON dataset, we use VITON-HD dataset for the unified evaluation. Re-training previous methods cannot guarantee the optimal parameters for fair comparison, so we use officially pretrained weights from each respective method for inference. Although the two datasets come from the same source which indicate their distributions are similar, we still mark the methods which are trained on VITON dataset while tested on VITON-HD dataset as cross-dataset evaluation.
The resolution of generated images follows the implementation of officially released code. Specifically, the resolution of VITON, CP-VTON, LM-VTON, CP-VTON+, PFAFN, DCTON, Style-Flow and SDAFN is $256\times 192$. The resolution of VITON-HD and HR-VITON is $1024\times 768$. The resolution of Ladi-VTON, DCI-VTON, GP-VTON, CAT-DM, Stable VITON and TPD is $512\times 384$. As the evaluation criteria are affected by image resolution, the methods with the same resolution are put together for comparison. CLIP [98] could take images with any resolution as input and resize the resolution to $224\times 224$ for later computation.
Figure 8: Two testing conditions are involved for evaluation. Paired evaluation refers to the condition that the clothing image and person image are paired and the clothing-agnostic image is parsed from the person image. Unpaired evaluation refers to that the clothes in the clothing image and person image are different, so there is no ground truth for the try-on result.